# Calibrating LLMs for Text-to-SQL Parsing by Leveraging Sub-clause Frequencies

### v1

@@ Abstract
Terrance Liuthanks: work was conducted during an internship at Bloomberg Carnegie Mellon University                        Shuyi Wang Bloomberg                        Daniel Preotiuc-Pietro Bloomberg
Yash Chandaranathanks: equal contribution. Bloomberg                        Chirag Guptafootnotemark: Bloomberg
@@ 3 Calibration
However, because $p$ can take on an infinite number of values between 00 and $1$, ensuring that Equation 1 holds for all $p$ is intractable (Gupta et al., 2020). A typical simplification is to consider calibration when the output of $s$ is discretized into coarser probability bins (e.g., [0.0.2),,[0.8.1.0])[0.0.2)$\ldots$,[0.8.1.0])[ 0.0.2 ) , … , [ 0.8.1.0 ] ) and ensuring calibration on those bins. This is the definition we will aim for.
@@ 4.2 Multivariate Platt Scaling
Although Equation 5 resembles typical ML models, we emphasize a key difference in that while signals $s_{i}$ are learned on the training data, the weights $\mathbf{w_{i}}$ are learned on a held out validation set—also called calibration data. Typically, the validation set is only used for model selection and not modeling itself so that we have some out-of-sample metric that checks for overfitting. In our case, since the secondary Platt scaling model is small ($m+1$ parameters), we can reuse validation data without significant overfitting, as long as a reasonable number of validation points are available (say $\Omega(m^{2})$). Multivariate Platt scaling is one of many “post-hoc" calibration approaches that aim to improve calibration of a model without sacrificing accuracy.
@@ 4.3 Extracting Sub-Clause Frequencies
Our method draws inspiration from self-consistency (Wang et al., 2022) for a model that produces multiple outputs for the same prompt. Specifically, each candidate output is assigned a frequency score that counts the number of times the candidate appears among all outputs. The one with the highest frequency score (i.e., the most consistently outputted candidate) is selected as the final output.
@@
As a result, we produce signals $s_{i}$ for every sub-clause, to which we apply multivariate Platt scaling as described in the previous subsection. In Figure 2, we provide a simplified example of what this procedure looks like using nucleus sampling (Holtzman et al., 2020) and beam search (Freitag and Al-Onaizan, 2017). Appendix B describes our procedure in full detail.
@@ 5.2 Models
We use T5-3B-PICARD (which we will refer to as T5 3B going forward) and Llama 3.1 8B Instruct on the SPIDER dataset. The T5-3B-PICARD model is a T5 model (available at this link) fine-tuned on SPIDER using constrained decoding for text-to-SQL parsing, achieving strong results on the task (Scholak et al., 2021). Llama 3.1 8B Instruct (Dubey et al., 2024) is used with zero-shot prompting because the model is also able to obtain strong accuracy on this task. For the BIRD dataset, we use both Llama 3.1 8B Instruct and Llama 3.1 70B Instruct (Dubey et al., 2024) with zero-shot prompting as they also have been shown to perform relatively well. We did not use T5 variants for this dataset as their accuracies were low ($23.34\%$) (Li et al., 2024). Prompts for our experiments are presented in Appendix C.
@@ 5.3 Computing Sub-Clause Frequencies
To run MPS using parsed sub-clause frequencies as signals, we sample outputs via nucleus sampling and beam search. In our experiments, we set the number of samples for each method to $k=10$, [^2] These numbers were set after initial experiments on the development set. We note that we did not attempt to further tune these hyperparameters, which can only lead to better results for our proposed approach. leading to a total of 20 samples. Beam search takes the top $k=10$ outputs by probability (after length normalization) using a beam width of $2k$. Then, we compute sub-clause frequencies for each sample using all samples from nucleus sampling and beam search (as illustrated in Figure 2).
We note that we remove a small proportion of outputs that have syntactical errors since the correctness of such outputs can be determined quite easily with automated SQL parsers.
@@ 5.4 Metrics for Calibration
    expected calibration error (ECE): bins $S_{i}$ are defined as $k$ equal-width bins: $[0,\frac{1}{k}),[\frac{1}{k},\frac{2}{k}),\ldots,[\frac{k-1}{k},1]$.
@@
    adaptive calibration error (ACE): bins $S_{i}$ are defined such that the all sizes of all bins are the same (i.e., $P_{\mathcal{X\sim D}}(s(X)\in S_{i})=\frac{1}{k}$).
In addition to minimizing calibration error, it is often desirable to output a wide range of scores $s$ that can help users better distinguish between high and low confidence outputs. The spread of scores is sometimes referred to as resolution (Bröcker, 2009) in the literature, where higher resolution corresponds to a larger spread. Calibration error, however, does not capture this spread. For example, as an extreme case, an algorithm could output the same probability $s(x)=P_{D}(Y=1)$ for all examples in $x\in D$. Despite being perfectly calibrated (i.e., calibration error of 00), this scoring function serves no purpose in communicating risk for individual examples $X$.
@@
Intuitively, a lower Brier score is desirable since correct outputs ($Y=1$) should be accompanied by higher probability estimates ($s(X)$ closer to $1$) and incorrect outputs ($Y=0$) by lower probability estimates ($s(X)$ closer to 00). Moreover, unlike calibration error, Brier score penalizes algorithms with low resolution. [^3] In fact, as shown in Bröcker (2009, Equation 13), Brier score can be decomposed into calibration error and (negative) resolution, thereby capturing both properties that are valuable for uncertainty quantification.
@@ 6.1 Calibration
Table 1 presents the results of our LLM-based text-to-SQL parsers on both datasets using the two approaches to calibration: Platt scaling (PS) and multivariate Platt scaling (MPS) with sub-clause frequencies (SCF). Here, MPS uses SCF scores derived from both nucleus sampling and beam search. As reference, we report results for uncalibrated probabilities (Stengel-Eskin and Van Durme, 2023) that are outputted directly from the model.
We observe that the probabilities produced by the LLMs are consistently and significantly miscalibrated across all experiments. However, Platt scaling (PS) improves calibration across both datasets and models. In some cases, PS improves calibration metrics by very wide margins (e.g. Llama 3.1 8B Instruct on SPIDER shows a reduction of $0.17$ in Brier score and $0.37$ in ECE and ACE).
@@ 7.1 Calibration Plots
Figure 3 shows the calibration plots using ECE, comparing probabilities calibrated by PS and MPS. To convey the number of examples in each bin, we scale the radius of each marker by the bin size. Generally, we observe that calibration errors are closer to 00 (i.e., points lie closer to $y=x$) for MPS compared to PS, reflecting the lower ECE of MPS.
@@ 7.2 Comparing Calibration across Models
When examining (uncalibrated) probabilities outputted directly by the LLM on SPIDER, we observe that those produced by fine-tuned T5 3B are better calibrated (i.e., Brier score, ECE, and ACE are almost $2$x smaller) than those of the larger Llama 3.1 8B Instruct model, despite the two models achieving similar test accuracies (Table 3). However, after applying (multivariate) Platt scaling, the differences become small, providing strong evidence that when evaluating calibration for LLMs, post-hoc calibration (e.g., Platt scaling) should be conducted before forming conclusions about how well-calibrated LLMs are relative to each other.
@@ 11 Acknowledgements
This work was supported by a Bloomberg Data Science Fellowship to the first author.
@@ Appendix A Supplementary experiments
Varying base uncertainty scores used for post-hoc calibration.
In Table 2, we ran Platt scaling (PS) on additional base uncertainty scores using outputs from Llama 3.1 8B Instruct on the SPIDER dataset. Specifically, we consider perplexity, P(True) (Kadavath et al., 2022), and verbalized confidence (Tian et al., 2023). While applying PS to verbalized confidence performs poorly, PS is able to achieve low calibration error on the other base scores. However Brier score and AUC vary greatly. For example, PS on perplexity performs poorly w.r.t to Brier score and AUC but the best w.r.t ECE and ACE. Finally, we stress that MPS using SCF scores (Table 1) still performs better than all these configurations of PS, highlighting the effectiveness of our proposed method.
Base score Brier ECE ACE AUC
Model Probabilities 0.173 0.044 0.045 0.712
Perplexity 0.185 0.022 0.028 0.613
P(True) 0.177 0.027 0.037 0.678
Verbalized Confidence 0.19 0.034 0.052 0.564
@@
In Table 3 we show model performance on SPIDER and BIRD. For our experiments, we sample $10$ outputs using nucleus sampling and $10$ using beam search. Then for each query, we take the model output (among all outputs sampled) with the highest model probability (sum over log token probabilities) as our prediction. Table 3 shows the performance when choosing outputs in this way, picking the top output among the set of samples outputted by (1) nucleus sampling (Nucleus), (2) beam search (Beam), and (3) both (N+B). Generally speaking, we do not see significant different in model accuracy among these three choices and thus choose (3) for our experiments out of convenience. We also note that that Llama 3.1 70B Instruct is a very strong model, performing on par or better on BIRD than various published methods that combine GPT-4 with more advanced prompting strategies (Pourreza and Rafiei, 2024; Gao et al., 2023).
@@
We see that across all strata, MPS produces probabilities that are more calibrated, suggesting that queries for which MPS and PS disagree on the most are those for which MPS is better calibrated on. For example, in the top 1%, PS on average produces a probability of $0.464$ while MPS produces an average probability of $0.825$, which is closer to the true proportion of correct samples of $0.714$. Full results are in the Table 5, where we show how and to what degree MPS shifts its calibrated probability output when compared to PS. In particular, we show the average probability for top $1\%$, $5\%$, $10\%$, and $20\%$ samples, sorted by the change $\Delta$ in probability between MPS and PS. We see that MPS shifts the probabilities so that for each set of examples, the calibration error is smaller when compared to PS. Finally, we show examples of queries with the largest absolute difference in probability between MPS and PS in the Tables 8, 9, 10 and 11.
@@
Note that while for completeness, we report Brier score in Tables 6 and 7, we cannot use Brier score to compare between levels of difficulty since the model accuracy for each difficulty varies substantially (See Section 7.2) for a more detailed explanation.
@@
In Tables 8, 9, 10, 11, we show specific examples of queries from the SPIDER and BIRD datasets for which MPS and PS differ the most. In particular, we show examples of correct outputs where MPS shifts the calibrated probability upwards the most and incorrect outputs where MPS shifts probabilities downwards.

### v2

@@ Abstract
Terrance Liuthanks: Work done during an internship at Bloomberg Carnegie Mellon University          Shuyi Wang Bloomberg          Daniel Preoţiuc-Pietro Bloomberg
Yash Chandaranathanks: equal contribution Bloomberg          Chirag Guptafootnotemark: Bloomberg
@@ 3 Calibration
However, because $p$ can take on an infinite number of values between 0 and 11, ensuring that Equation 1 holds for all $p$ is intractable (Gupta et al., 2020). A typical simplification is to consider calibration when the output of $s$ is discretized into coarser probability bins (e.g., $0.0.2),\ldots,[0.8.1.0])$ and ensuring calibration on those bins. This is the definition we will aim for.
@@ 4.2 Multivariate Platt Scaling
Although Equation 5 resembles typical ML models, we emphasize a key difference in that while signals $s_{i}$ are learned on the training data, the weights $\mathbf{w_{i}}$ are learned on a held out validation set—also called calibration data. Typically, the validation set is only used for model selection and not modeling itself so that we have some out-of-sample metric that checks for overfitting. In our case, since the secondary Platt scaling model is small ($m+1$ parameters), we can reuse validation data without significant overfitting, as long as a reasonable number of validation points are available.222say $\approx 10m$, which is based on the sample complexity of logistic regression (see Hsu and Mazumdar (2024)). Multivariate Platt scaling is one of many “post-hoc" calibration approaches that aim to improve calibration of a model without sacrificing accuracy.
@@ 4.3 Extracting Sub-Clause Frequencies
Our method draws inspiration from self-consistency (Wang et al., 2022) for a model that produces multiple outputs for the same prompt. Specifically, each candidate output is assigned a frequency score that counts the number of times the candidate appears among all outputs. The one with the highest frequency score (i.e., the most consistently generated candidate) is selected as the final output.
@@
As a result, we produce signals $s_{i}$ for every sub-clause, to which we apply multivariate Platt scaling as described in the previous subsection. In Figure 2, we provide a simplified example of what this procedure looks like using nucleus sampling (Holtzman et al., 2020) and beam search (Freitag and Al-Onaizan, 2017). Appendix LABEL:appx:additional_scf describes our procedure in full detail.
@@ 5.2 Models
We use T5-3B-PICARD (which we will refer to as T5 3B going forward) and Llama 3.1 8B Instruct on the SPIDER dataset. The T5-3B-PICARD model is a T5 model (available at this link) fine-tuned on SPIDER using constrained decoding for text-to-SQL parsing, achieving strong results on the task (Scholak et al., 2021). Llama 3.1 8B Instruct (Dubey et al., 2024) is used with zero-shot prompting because the model is also able to obtain strong accuracy on this task. For the BIRD dataset, we use both Llama 3.1 8B Instruct and Llama 3.1 70B Instruct (Dubey et al., 2024) with zero-shot prompting as they also have been shown to perform relatively well. We did not use T5 variants for this dataset as their accuracies were low ($23.34\%$) (Li et al., 2024). Prompts for our experiments are presented in Appendix LABEL:appx:additional_exp_details.
@@ 5.3 Computing Sub-Clause Frequencies
To run MPS using parsed sub-clause frequencies as signals, we sample outputs via nucleus sampling and beam search. In our experiments, we set the number of samples for each method to k=10k=10, [^3] These numbers were set after initial experiments on the development set. We note that we did not attempt to further tune these hyperparameters, which can only lead to better results for our proposed approach. leading to a total of 20 samples. Beam search takes the top $k=10$ outputs by probability (after length normalization) using a beam width of $2k$. Then, we compute sub-clause frequencies for each sample using all samples from nucleus sampling and beam search (as illustrated in Figure 2).
We note that we remove a small proportion of outputs that have syntactic errors since the correctness of such outputs can be determined quite easily with automated SQL parsers.
@@ 5.4 Metrics for Calibration
    expected calibration error (ECE) (Naeini et al., 2015): bins $S_{i}$ are defined as $k$ equal-width bins: $0,\frac{1}{k}),[\frac{1}{k},\frac{2}{k}),\ldots,[\frac{k-1}{k},1]$.
@@
    adaptive calibration error (ACE) (Nixon et al., 2019): bins $S_{i}$ are defined such that the all sizes of all bins are the same (i.e., $P_{\mathcal{X\sim D}}(s(X)\in S_{i})=\frac{1}{k}$).
In addition to minimizing calibration error, it is often desirable to output a wide range of scores $s$ that can help users better distinguish between high and low confidence outputs. The spread of scores is sometimes referred to as resolution (Bröcker, 2009) in the literature, where higher resolution corresponds to a larger spread. Calibration error, however, does not capture this spread. For example, as an extreme case, an algorithm could output the same probability $s(x)=P_{D}(Y=1)$ for all examples in $x\in D$. Despite being perfectly calibrated (i.e., calibration error of 0), this scoring function serves no purpose in communicating risk for individual examples $X$.
@@
Intuitively, a lower Brier score is desirable since correct outputs ($Y=1$) should be accompanied by higher probability estimates ($s(X)$ closer to $1$) and incorrect outputs ($Y=0$) by lower probability estimates ($s(X)$ closer to 0). Moreover, unlike calibration error, Brier score penalizes algorithms with low resolution. [^4] In fact, as shown in Bröcker (2009, Equation 13), Brier score can be decomposed into calibration error and (negative) resolution, thereby capturing both properties that are valuable for uncertainty quantification.
@@ 6.1 Calibration
Table 1 presents the results of our LLM-based text-to-SQL parsers on both datasets using the two approaches to calibration: Platt scaling (PS) and multivariate Platt scaling (MPS) with sub-clause frequencies (SCF). Here, MPS uses SCF scores derived from both nucleus sampling and beam search. As reference, we report results for uncalibrated probabilities (Stengel-Eskin and Van Durme, 2023) that are generated directly from the model.
We observe that the probabilities produced by the LLMs are consistently and significantly miscalibrated across all experiments. However, Platt scaling (PS) improves calibration across both datasets and models. In some cases, PS improves calibration metrics by very wide margins (e.g., Llama 3.1 8B Instruct on SPIDER shows a reduction of $0.17$ in Brier score and $0.37$ in ECE and ACE).
@@ 7.1 Calibration Plots
Figure 3 shows the calibration plots using ECE, comparing probabilities calibrated by PS and MPS. To convey the number of examples in each bin, we scale the radius of each marker by the bin size. Generally, we observe that calibration errors are closer to 0 (i.e., points lie closer to $y=x$) for MPS compared to PS, reflecting the lower ECE of MPS.
@@ 7.2 Comparing Calibration across Models
When examining (uncalibrated) probabilities generated directly by the LLM on SPIDER, we observe that those produced by fine-tuned T5 3B are better calibrated (i.e., Brier score, ECE, and ACE are almost 22x smaller) than those of the larger Llama 3.1 8B Instruct model, despite the two models achieving similar test accuracies (Table 5). However, after applying (multivariate) Platt scaling, the differences become small, providing strong evidence that when evaluating calibration for LLMs, post-hoc calibration (e.g., Platt scaling) should be conducted before forming conclusions about how well-calibrated LLMs are relative to each other.
@@ 7.3 Ablation Study for Deriving SCF Scores
7.4 Sensitivity analysis for MPS
In Table 2, we show the performance of MPS with the size of the calibration set. We see that the performance of MPS declines as the size of the calibration set decreases.
Next, we study the impact of the number of model outputs $k$ used to calculate SCF scores for MPS in 3. As expected, the larger number of outputs used increases performance of calibration, with a noticeable increase for more than 1 output and with $k=5$ as used in the experiments in this paper being a good trade-off between latency and performance and larger values bringing additional improvements.
dataset factor Brier ECE ACE AUC
SPIDER all 0.164 0.023 0.023 0.746
$2^{-1}$ 0.166 0.033 0.036 0.741
$2^{-2}$ 0.168 0.036 0.039 0.737
$2^{-3}$ 0.173 0.047 0.050 0.725
$2^{-4}$ 0.175 0.045 0.048 0.713
$2^{-5}$ 0.180 0.060 0.063 0.691
$2^{-6}$ 0.192 0.094 0.099 0.637
BIRD all 0.194 0.035 0.032 0.737
$2^{-1}$ 0.197 0.035 0.036 0.727
$2^{-2}$ 0.200 0.042 0.043 0.717
$2^{-3}$ 0.206 0.056 0.055 0.701
$2^{-4}$ 0.212 0.065 0.066 0.686
$2^{-5}$ 0.227 0.106 0.106 0.668
$2^{-6}$ 0.243 0.141 0.141 0.627
Table 2: We evaluate MPS on outputs from Llama 3.1 8B Instruct on SPIDER and BIRD when reducing the size of the calibration set. We sample a factor of the original calibration set to run MPS and evaluate on the test set, taking an average over 10 random runs.
Dataset k Brier ECE ACE AUC
SPIDER 10 0.164 0.023 0.023 0.746
5 0.167 0.027 0.030 0.732
3 0.168 0.028 0.025 0.726
1 0.170 0.040 0.038 0.717
BIRD 10 0.194 0.035 0.032 0.737
5 0.194 0.041 0.035 0.732
3 0.195 0.046 0.041 0.728
1 0.203 0.029 0.041 0.695
Table 3: We evaluate MPS on outputs from Llama 3.1 8B Instruct on SPIDER and BIRD when reducing the number of additional model outputs $k$ that are used to calculate SCF scores.
7.5 Uncertainty scores used for post-hoc calibration
We study additional methods for computing the base uncertainty scored used for calibration and consider specifically perplexity, P(True) (Kadavath et al., 2022), and verbalized confidence (Tian et al., 2023). We ran Platt scaling (PS) on additional base uncertainty scores using outputs from Llama 3.1 8B Instruct on the SPIDER dataset and present the results in Table 4.
While applying PS to verbalized confidence performs poorly, PS is able to achieve low calibration error on the other base scores. However Brier score and AUC vary greatly. For example, PS on perplexity performs poorly w.r.t to Brier score and AUC but the best w.r.t ECE and ACE. Finally, we stress that MPS using SCF scores (Table 1) still performs better than all these configurations of PS, highlighting the effectiveness of our proposed method.
Base score Brier ECE ACE AUC
Model Probabilities 0.173 0.044 0.045 0.712
Perplexity 0.185 0.022 0.028 0.613
P(True) 0.177 0.027 0.037 0.678
Verbalized Confidence 0.19 0.034 0.052 0.564
@@ 11 Acknowledgements
This research is supported by a grant from the Bloomberg Data Science Ph.D. Fellowship Program to the first author.
@@ Appendix A Supplementary experiments
In Table 5 we show model performance on SPIDER and BIRD. For our experiments, we sample $10$ outputs using nucleus sampling and $10$ using beam search. Then for each query, we take the model output (among all outputs sampled) with the highest model probability (sum over log token probabilities) as our prediction. Table 5 shows the performance when choosing outputs in this way, picking the top output among the set of samples generated by (1) nucleus sampling (Nucleus), (2) beam search (Beam), and (3) both (N+B). Generally speaking, we do not see significant different in model accuracy among these three choices and thus choose (3) for our experiments out of convenience. We also note that that Llama 3.1 70B Instruct is a very strong model, performing on par or better on BIRD than various published methods that combine GPT-4 with more advanced prompting strategies (Pourreza and Rafiei, 2024; Gao et al., 2023).
@@
We see that across all strata, MPS produces probabilities that are more calibrated, suggesting that queries for which MPS and PS disagree on the most are those for which MPS is better calibrated on. For example, in the top 1%, PS on average produces a probability of $0.464$ while MPS produces an average probability of $0.825$, which is closer to the true proportion of correct samples of $0.714$. Full results are in the Table 7, where we show how and to what degree MPS shifts its calibrated probability output when compared to PS. In particular, we show the average probability for top $1\%$, $5\%$, $10\%$, and $20\%$ samples, sorted by the change $\Delta$ in probability between MPS and PS. We see that MPS shifts the probabilities so that for each set of examples, the calibration error is smaller when compared to PS. Finally, we show examples of queries with the largest absolute difference in probability between MPS and PS in the Tables A, LABEL:tab:cal_ex_spider_llama_8b, LABEL:tab:cal_ex_bird_llama_8b and LABEL:tab:cal_ex_bird_llama_70b.
@@
Note that while for completeness, we report Brier score in Tables 8 and 9, we cannot use Brier score to compare between levels of difficulty since the model accuracy for each difficulty varies substantially (See Section 7.2) for a more detailed explanation.
Weight / Sub-clause Analysis of MPS.
We conduct additional analysis on the weights produced for MPS on outputs from Llama 3.1 8B Instruct. In Tables 10 and 11, we report the sum and absolute sum of (standardized) weights corresponding to each sub-clause type for SPIDER and BIRD respectively. Generally, the learned feature weights represent to what extent the final probability of correctness should be shifted, given how consistent the language model is in generating a particular sub-clause. In addition, conduct an ablation study in which we remove features corresponding to each sub-clause type as input into MPS. Tables 12 and 13 show the impact of each sub-clause on the calibration metrics for SPIDER and BIRD respectively.
We find that our aggregate sub-clause score (AGG) is the most significant feature utilized by MPS, which is expected given that it aggregates information from all sub-clauses (i.e., product of SCF scores). Not only does it have the highest weight (Tables 10 and 11), but removing it from MPS most significantly degrades ACE and ECE for both SPIDER and BIRD (Tables 12 and 13). On the other hand, we find that the SCF scores corresponding to DISTINCT, HAVING, LIMIT, and set operations have small weights (Tables 10 and 11). Removing these clauses also has minimal impact on MPS’s performance (Tables 12 and 13). While trends for the remaining sub-clauses are not as consistent between the two datasets, they are still unsurprising. For example, removing SELECT and WHERE features, two of the more important and common sub-clauses for SQL, significantly impacts MPS for BIRD, especially with respect to AUC. Finally, we note that while the improvement of MPS over PS with respect to Brier score illustrates that incorporating all sub-clauses as a whole is important for calibration (i.e., our main empirical result), Tables 12 and 13 show that removing any individual sub-clause does not significantly impact Brier score for MPS.
sub-clause abs. weight ($\downarrow$) weight
FROM 0.826 0.227
AGG 0.687 0.687
SELECT 0.507 -0.137
ORDER 0.282 0.134
WHERE 0.260 0.082
GROUP 0.199 0.199
ON 0.191 -0.186
HAVING 0.168 0.129
DISTINCT 0.131 0.131
LIMIT 0.102 0.102
Set Op. 0.038 0.038
Table 10: We report the sum and absolute sum of (standardized) weights corresponding to each sub-clause type for MPS trained on Llama 3.1 8B Instruct outputs on SPIDER.
sub-clause abs. weight ($\downarrow$) weight
AGG 0.343 0.343
SELECT 0.310 0.302
WHERE 0.255 0.242
ORDER 0.224 0.087
ON 0.205 -0.158
FROM 0.180 0.172
GROUP 0.178 0.165
DISTINCT 0.081 -0.010
HAVING 0.077 0.041
LIMIT 0.066 -0.022
Set Op. 0.029 0.016
Table 11: We report the sum and absolute sum of (standardized) weights corresponding to each sub-clause type for MPS trained on Llama 3.1 8B Instruct outputs on BIRD.
Removed Feature Brier ECE ACE AUC
0.164 0.023 0.023 0.746
AGG 0.001 0.006 0.008 0.000
Set Op. 0.000 0.000 0.000 -0.000
DISTINCT -0.000 -0.002 0.004 0.001
SELECT -0.001 -0.001 0.005 0.002
FROM 0.001 0.002 0.002 -0.005
ON 0.000 0.000 0.000 0.001
WHERE -0.000 0.005 0.002 0.002
GROUP -0.001 0.006 0.008 0.004
HAVING 0.000 0.000 0.000 -0.001
ORDER 0.000 -0.000 -0.000 -0.000
LIMIT -0.000 0.001 0.001 0.000
Table 12: We report the change in performance of MPS when sub-clauses are removed as input signals. The ablation study is conducted on MPS trained using Llama 3.1 8B Instruct outputs for SPIDER. We bold values that are significantly large (delta greater than 0.005).
Removed Feature Brier ECE ACE AUC
0.194 0.035 0.032 0.737
AGG 0.001 0.016 0.015 -0.003
Set Op. -0.000 -0.000 -0.000 0.000
DISTINCT -0.000 0.002 0.001 0.001
SELECT 0.005 0.002 0.008 -0.017
FROM 0.000 -0.001 0.003 0.000
ON 0.001 0.005 0.000 -0.003
WHERE 0.002 -0.000 0.008 -0.007
GROUP -0.000 0.005 -0.000 0.000
HAVING -0.000 0.002 0.000 0.000
ORDER 0.000 0.006 0.005 0.001
LIMIT 0.000 0.001 0.001 -0.001
Table 13: We report the change in performance of MPS when sub-clauses are removed as input signals. The ablation study is conducted on MPS trained using Llama 3.1 8B Instruct outputs for BIRD. We bold values that are significantly large (delta greater than 0.005).
@@
In Tables A, LABEL:tab:cal_ex_spider_llama_8b, LABEL:tab:cal_ex_bird_llama_8b, LABEL:tab:cal_ex_bird_llama_70b, we show specific examples of queries from the SPIDER and BIRD datasets for which MPS and PS differ the most. In particular, we show examples of correct outputs where MPS shifts the calibrated probability upwards the most and incorrect outputs where MPS shifts probabilities downwards.